OpenStudy (anonymous):

i need some help with some college algebra. If I have to sell pizza for $15 and the cost of making the pizzas include a fixed cost of $55 and a labor cost of $4. What would the equation be to determine the revenue. Determine the total cost. How many pizzas must be sold to break even.

OpenStudy (anonymous):

Your wording is confusing me. Is that a labor cost of $4 per pizza? A fixed cost of $55 per pizza?

OpenStudy (anonymous):

you are going to lose money on this venture

OpenStudy (anonymous):

oh i see you need to solve \[55+4x=15x\] for x

OpenStudy (anonymous):

you get \[55=11x\] \[x=5\]

OpenStudy (anonymous):

Cost C (q) = 4q + 55 Revenue R(q) = 15q -> Break even: 15q = 4q + 55 => 11q = 55 Thus q = 55/11 = 5 pizza In order to break even, the quantity sale out need to be at least 5 pizzas.